A tall vertical cylinder filled with water is kept on a horizontal table top. Two small holes A and B are made on the wall of the cylinder, A near the middle and B just below the free surface of water. State and explain your observation

Grade:9

2 Answers

Susmita
5 years ago
Let the column be of height H.
Let h1 and h2 be the distance of B and A from the top(or free surface) of water column.
So as B is just below the free surface so h1 is very small or almost zero.
As A is in the middle of the column so h2=H/2.
Then we know velocity of efflux
v=root(2gh)
Let v1 is the velocity of liquid coming out through hole B.So v1=root(2gh1) which  very small or almost zero.
If v2 is the velocity of efflux through A then
v2=root(2gh2)=root(2gH/2).This is obviously greater than v1.This is our 1st observation.
Our second observation is the range is maximum for efflux through A as h=H/2.Means water coming out through A will fall far away from the cylinder than water from B.
If it helps please approve.
 
Manvi jain
3 years ago
The water jet through A will have a larger horizontal range because the water pressure is more at that point. While B is just below the surface so water will come out of it without any significant pressure or jet. It will just leak through it